A ball is thrown with velocity v it reaches maximum height.if one wishes to triple the maximum height then the ball should be thrown with what velocity

Grade:11

4 Answers

Vikas TU
6 years ago
We definitely realize that , 
Height(h) = u2/2g 
In any case, its speed is v 
At that point Height (h) = v2/2g 
    v = starting speed 
    g = increasing speed because of gravity 
    h2/h1 = v1 2/v2 2 
    √( h2/h1) = v1/v2 
    v2√( h2/h1) = v1 
    h2 = 3 h1 
    v2√3 h1/h1 = v1 
    v1 = √3 v2 
User
6 years ago
Root3v must be the answer because according to question it is given that the when the height is tripled than velocity is unknown but we know rh
gavini haritha
6 years ago
We know that h is directly proportional to u^2. Then h/3h=u^2/x (where x the velocity when height is 3h),by this we get x=3u^2
Omed prakash
6 years ago
H1= Vo2/2g and 3H1= v2/2gAgain divided both equation H1/3H1=Vo2/v2 Vo/v= squareroot of H1/3H1Then v= square root of 3 Vo